OpenStudy (anonymous):

PLEASE HELP me in algebra 2 lesson 8 unit 6

OpenStudy (anonymous):

What r the questions??

OpenStudy (anonymous):

what is the solution to the equation 

OpenStudy (anonymous):

2/13 1/12 -8/5 13/2

OpenStudy (anonymous):

You must have written down the problem incorrectly, none of these are the correct answers.

OpenStudy (anonymous):

it was suppose to be (-9x+5) sorry

OpenStudy (anonymous):

Start out by squaring both sides, to extinguish the 1/2 (note: raising something to (1/n) is taking the nth root of that something; ie a^(1/2) means you're taking the square root of a) $$-9x+5 = 3+4x $$ Now let's isolate x. We'll carry -9x over from the left, and add it to the right. $$5 = 3+13x$$ Now carry +3 from the right, and subtract it from the left. $$2 = 13x$$ Finally, divide the left side by 13 $$2/13 = x$$

OpenStudy (anonymous):

ok thats what i got but i wasnt sure! Let f(x) = –2x + 4 and g(x) = –6x – 7. Find f(x) – g(x). (2 points)

OpenStudy (anonymous):

Alright, for this one, it's simple substitution. $$[f(x) - g(x)] \to [(-2x + 4) - (-6x - 7)]$$ Distribute the negative into the right hand function, and now we can get rid of the parenthesis. $$-2x + 4 + 6x + 7$$ And finally we have, $$ 4x +11 $$

OpenStudy (anonymous):

thank you for helping me! What is the inverse of the given relation? y = 3x + 12 (2 points)

OpenStudy (anonymous):

Alright, for this one we want to solve for x. We can firstly subtract 12 from both sides, which will give us what?

OpenStudy (anonymous):

y-12=3x

OpenStudy (anonymous):

Right. Now we want to isolate x (just like I did in the first problem). How can we do that from here?

OpenStudy (anonymous):

divide each by 3 and get y-4=x

OpenStudy (anonymous):

so thats the answer?

OpenStudy (anonymous):

Almost! You need to divide the entire left side by 3. So that it becomes $$\frac{y-12}{3} = x$$

OpenStudy (anonymous):

Now here comes the final step. Simple swap x & y. Nothing special, just swap places. $$ \frac{x-12}{3} = y $$

OpenStudy (anonymous):

And that is your inverse of the given relation.

OpenStudy (anonymous):

oh okay!